Two blocks (2.8 and 1.9kg) are connected by a light string that passes over a pulley. The top block is connected to a spring, which is initially unstretched. The spring has a spring constant of 110N/m and is capable of stretching great lengths. Once the 1.9kg block is released, how far will it descend before momentarily stopping?
Answer #1

here,

mass of block 1, m1 = 2.8 kg

mass of bloxk 2 , m2 = 1.9 kg

spring constant , K = 110 N/m

let the block 1.9 kg travels x m before stopping

equating the forces

as the net force on the sysetm is zero when it stops

(m1 - m2) * g = K * x

(2.8 - 1.9) * 9.81 = 110 * x

x = 0.0803 m = 8.03 cm

the block travels 8.03 cm before stoping